OpenStudy (pphalke):

Sarah Meeham blends coffee for Tasti-Delight. She needs to prepare 160 pounds of blended coffee beans selling for $3.75 per pound. She plans to do this by blending together a high quality bean costing $6.00 per pound and a cheaper bean at $2.00 per pound. To the nearest pound, find how much high quality coffee bean and how much cheaper bean she should blend

OpenStudy (john_es):

First you should identify your variables. Let's x be the quantity of cheap coffe and y quantity of expensive coffe. How much quantity of coffe (of both types) the problem says you need?

OpenStudy (pphalke):

so i would do x and y first but how would i make an equation

OpenStudy (smartnerd1111):

Call the quantity of high-quality coffee is X Cheaper coffee bean is Y Since the total amount of coffee is 149 pound so Have: X+Y= 149

OpenStudy (john_es):

Ok, the problem says you need 160 pounds of coffe, right?

OpenStudy (john_es):

Yeah, this is the first equation.

OpenStudy (john_es):

Well, but with 160 instead 149.

OpenStudy (smartnerd1111):

oh

OpenStudy (john_es):

So the total amount of coffe you need is, \[x+y=160\] Ok?

OpenStudy (pphalke):

okay

OpenStudy (john_es):

Well, now that you have the equation for the amount of coffe, you need the equation for the amount of money. The total money for the cheap coffee is: \[2\cdot x \] The total money for the expensive coffee is \[6\cdot y\] The total money you will have is, \[3.75\cdot160=600\] So the equation is, \[2x+6y=600\]

OpenStudy (john_es):

Now you have your system of equations, \[x+y=160\\ 2x+6y=600\] Do you know how to solve it?

OpenStudy (pphalke):

not really i knowu have to get y alone right?

OpenStudy (john_es):

Yes, you can do it by substituion. In the first equation, solve for y you will get, \[y=160-x\] Ok?

OpenStudy (john_es):

Now, plug this into the second equation, you will get, \[2x+6(160-x)=600\]Ok? Now solve for x

OpenStudy (john_es):

You will get something like this, \[2x+960-6x=600\Rightarrow -4x=-360\Rightarrow x=90\]

OpenStudy (john_es):

Now, with this value you can take whatever equation you want, plug in the value of x and obtain the value of y, for example, for the first equation \[x+y=160\Rightarrow 90+y=160\Rightarrow y=70\]. And that is your solution \[x=90, y=70\]

OpenStudy (john_es):

Do you understand it?

OpenStudy (pphalke):

yes

OpenStudy (john_es):

Ok, perfect. Just if you need more problems like, remember that all of them solve the same way.